OpenStudy (anonymous):

Solve for f. seven-twelfths times f equals three-fifths Thirty-five thirty-sixths One and thirty-five thirty-sixths Twenty-one sixtieths One and one thirty-fifth

OpenStudy (anonymous):

Okay so....

OpenStudy (anonymous):

\[ \frac{7}{12}f=\frac{3}{5} \]

OpenStudy (anonymous):

yeah and you do 7*12/3 right @wio - 5

OpenStudy (anonymous):

I don't know how to read that.

OpenStudy (anonymous):

Dividing by a fraction is multiplying by the reciprocal.\[ f = \frac{3}{5} \cdot \frac{12}{7} \]

OpenStudy (anonymous):

What is MLP and what is 312856?

OpenStudy (anonymous):

nothing Just a number y

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

did you get the answer?

OpenStudy (anonymous):

no @wio

OpenStudy (anonymous):

Simplify:\[ f = \frac{3}{5} \cdot \frac{12}{7} \]

OpenStudy (anonymous):

36 over 35 is that right @wio

OpenStudy (anonymous):

or is it 1 and 1/35

OpenStudy (anonymous):

@wio are you still here

OpenStudy (anonymous):

Both are correct.